### v12

# D4C Glove-train: Solving the Visual Abstract Reasoning Problems by Circumscribing and Building Distribution for Concepts

Abstract
This paper achieves noteworthy progress in the realm of abstract reasoning, particularly in addressing Raven’s Progressive Matrices (RPM) and Bongard-Logo challenges. Initially, we introduce Lico-Net, a novel baseline model that resolves RPM problems with remarkable accuracy. Leveraging this foundation, we advance with the D3C approach, which advocates representing the underlying concepts in abstract reasoning problems through distributions. This perspective enhances the performance of both Lico-Net and a baseline model excelling in Bongard-Logo tasks. To bolster the computational efficiency of D3C, we present the D3C-cos variant, offering a streamlined solution. Furthermore, we propose the D2C method, redefining concept boundaries within these domains and bridging the divide between high-level abstractions and their lower-dimensional counterparts. Finally, we extend our methodology to D4C, employing adversarial techniques to refine concept boundaries further and demonstrate substantial improvements in both RPM and Bongard-Logo challenges. Overall, our contributions present a fresh outlook and practical advancements in the field of abstract reasoning.
Index Terms:
Abstract reasoning, Raven’s Progressive Matrices, Bongard-logo problem, Adversarial Learning.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Deep neural networks have made remarkable achievements in various fields, including computer vision[1, 2, 3], natural language processing[4, 5, 6], generative model[7, 8, 9], visual question answering[10, 11], and visual abstract reasoning[14, 12, 13]. However, deep learning faces challenges in visual abstract reasoning, as it requires models to capture complex patterns, handle compositionality, and possess abstraction capabilities. Models must also be flexible and scalable to adapt to diverse graphical structures. Existing deep learning models often lack the necessary architectural rigor and representational power to address these challenges adequately. Overcoming these obstacles demands innovative model designs, enhanced representation techniques, and efficient algorithms. Progress in this area can lead to significant advancements in graphical reasoning and practical solutions.
Due to the higher demands of reasoning tasks compared to recognition, perception, and regression tasks, numerous prestigious research institutions are consecutively dedicating their efforts to studying abstract reasoning problems. They are focused on exploring diverse methodologies to address these challenges from multiple perspectives, with the expectation that the techniques developed in this field can be applied to benefit other areas of artificial intelligence. Notably, Ravens Progressive Matrices (RPM) challenges[14] and Bongard problems[12, 13] involve learning that ranges from basic perception to complex reasoning.
I-A RAVEN and PGM
The RAVEN database[15]—a unique compilation of Ravens Progressive Matrices (RPM) problems[14]—stands out due to its distinctive design elements. In each RAVEN problem, sixteen images are presented, with eight images forming the core of the challenge, referred to as the “statement.” The remaining eight images serve as a selection pool for subjects to choose from. The task requires selecting appropriate images from this pool to create, alongside the stem images, a 3x3 matrix of graphical patterns. This matrix showcases a progressive geometric pattern along the rows, encapsulating an abstract concept known as the “rule.” Figure 1 illustrates a specific example of a RAVEN case, providing insight into the methodology and general characteristics of these problems. Akin to RAVEN, PGM utilizes a format featuring eight images for the stem and answer pool. Nevertheless, the “rule” concept in PGM encompasses both row and column progressions. Figure 2 showcases a representative example of a PGM case.
Figure 1: RAVEN case
Figure 2: PGM case
I-B Bongrad-logo
The Bongard problem[12, 13] stands in stark contrast to the RPM problem, falling under the umbrella of small sample learning problems. Typically, Bongard problems consist of multiple images divided into two distinct groups: primary and auxiliary. Images within the primary group all convey an abstract concept that is bound by a specific rule, whereas images in the auxiliary group deviate from these rules in varying degrees. The crux of solving Bongard problems lies in the ability of deep learning algorithms to accurately classify ungrouped images into their respective categories. Notably, Bongard-logo[13] problems, a concrete instantiation of Bongard problems within the realm of abstract reasoning and a combination of a lot of Bongard-logo cases, pose a formidable challenge in terms of inferential complexity. Each Bongard-logo case comprises 14 images, including six images from both the primary and auxiliary groups, along with two images awaiting classification. The grouping criterion primarily relies on the geometric shapes present within the images and their arrangement. Figure 3 illustrates a representative Bongard-logo case, where every Bongard case consists of two sets of images: Set A (primary) and Set B (auxiliary). Set A contains six images, with geometric entities in each image adhering to a particular set of rules, while Set B encompasses six images that do not conform to the rules observed in Set A. The task entails assessing whether the images in the test set align with the rules manifested in Set A. The level of difficulty varies depending on the case’s structural complexity.
Figure 3: Bongard-Logo case
Bongard-logo problems can be categorized into three parts based on concept domains: 1) Free-form problems (FF), featuring shapes made of random action strokes; 2) Basic shape problems (BA), recognizing pattern of shape combinations; 3) High-level concept problems (HD), assessing abstract concept reasoning, like convexity and symmetry.
II Related work
II-A RPM solver
Discriminative models for image reasoning, such as CoPINet[19], LEN+teacher[18], and DCNet[20], offer diverse solutions with multi-dimensional outputs, focusing on learning differences, latent rules, and comparing disparities. NCD[22], SCL[23], SAVIR-T[24], and neural symbolism systems like PrAE, NVSA and ALANS[25, 26, 27] incorporate various techniques to improve reasoning accuracy and interpretability. By combining effective methods, RS-CNN and RS-TRAN[28] have achieved impressive results on the RPM problem. Triple-CFN implicitly extracts information on concepts and reasoning units, and “indexes” these reasoning unit information according to the concepts, achieving noteworthy reasoning accuracy. CRAB[29], based on a Bayesian modeling approach, has established a highly customized “conservatory” that aligns seamlessly with its proprietary methodology, embodied in the creation of a new RAVEN database. This bespoke conservatory, despite forfeiting the inherent core challenges of RAVEN such as solution diversity and uncertainty, has facilitated remarkable breakthroughs for CRAB. Meanwhile, the scientific community eagerly anticipates the profound implications of this innovative approach for future research endeavors.
II-B Bongard solver
Researchers have tackled Bongard problems via three primary methods: language feature models, convolutional neural networks, and generated datasets. Depweg et al. used a formal language and symbolic vocabulary with Bayesian inference, but this struggled with complex problems and had scalability issues[12]. Kharagorgiev and Yun pre-trained on image datasets for feature extraction[30]. Nie et al. experimented with CNNs and meta-learning but had subpar results[13]. These methods have distinct strengths and weaknesses, underscoring the need for more research. Yuan introduced PMoC[31], a probabilistic model that measures the likelihood of auxiliary group samples fitting the main group’s distribution, a novel and effective approach. Notably, Triple-CFN[32] has successfully applied a uniform approach to both Bongard-Logo and RPM problems.
II-C Distribution distance measurement method
II-C1 Wasserstein distance
The formula for the Wasserstein distance[33] can be expressed as:
$W(P,Q)=\mathbf{inf}(\sum|\,|\,X-Y|\,|\,)$ (1)
where $P$ and $Q$ are two probability distributions, and $X$ and $Y$ are random variables sampled from $P$ and $Q$, respectively. Here, $|\,|\,X-Y|\,|\,$ represents the distance between $X$ and $Y$, which can be the Euclidean distance, Manhattan distance, or other suitable metrics. The infimum (inf) denotes the minimum value among all possible mapping strategies. The Wasserstein distance is a measure of dissimilarity between distributions that accounts for the degree of difference between them, particularly suitable for handling non-uniform distributions and considering the shape and structure of the distributions. In the field of machine learning, the Wasserstein distance is widely used in generative models to evaluate the discrepancy between images generated by the generator and real images.
II-C2 Sinkhorn distance
Sinkhorn distance[34] is a metric based on optimal transport, which approximates the Wasserstein distance through the introduction of entropy regularization, leading to increased computational efficiency.
II-D Generative models and spectral normalization
Wasserstein GAN[36] is a deep learning model that uses the Wasserstein distance as its loss function in Generative Adversarial Networks (GANs) to improve training stability and sample quality.
Spectral normalization[35] is a stabilization technique for deep neural networks, particularly beneficial for GANs. It controls the Lipschitz constant by normalizing weight matrices, ensuring outputs change proportionally to inputs. This regularizes the network, preventing overfitting and instability. In GANs, it stabilizes adversarial training, improving convergence and sample quality. Spectral normalization is computationally efficient and offers a direct way to control network complexity, making it a valuable tool for deep learning practitioners.
III Methodology
Both Bongard-logo and RPM problems[14] belong to the category of abstract reasoning problems, which require solvers to possess the ability to extract specific concepts from abstract entities. These specific concepts, compared to lower-level concepts such as “pixel arrangement and configuration styles”, possess a higher level of abstraction and incorporate prior knowledge from human perspectives to varying degrees, encompassing aspects such as shape, size, color, positional relationships, graph concavity and convexity, and graph closure. Bongard-logo and RPM problems employ different questioning approaches aimed at evaluating the solver’s proficiency in extracting and learning these abstract concepts.
This paper presents Lico-Net, a novel baseline specifically designed for the RPM problem, which achieves remarkable performance in addressing this cognitive challenge. Subsequently, we introduce a fundamental argument that representing abstract and complex concepts as distributions, rather than vectors, is a more suitable approach. Based on this insight, we develop the D3C method. This technique not only enhances Lico-Net’s performance and interpretability on RPM tasks but also elevates the reasoning accuracy of Triple-CFN, a notable deep model that has excelled in Bongard-Logo problems. Recognizing the sources of reasoning difficulties in Bongard-Logo problem, we devise the D2C method. This method is broadly applicable to both RPM and Bongard-Logo challenges, demonstrating improved performance over several baseline models in these domains. Finally, we introduce the D4C method, which targets the essential aspects of abstract reasoning. This approach contributes significantly to advancing our understanding of RPM and Bongard-Logo problems, paving the way for future research in abstract reasoning.
IV Discriminator of Case (DC)
In this section, we introduce Lico-Net, a new baseline architecture designed to solve the RPM problem more effectively. This network utilizes advanced techniques for superior performance. Our evaluation shows Lico-Net’s effectiveness in RPM, opening new avenues for future research. Additionally, we critically assess the existing baseline for the Bongard-Logo problem, exploring its strengths and weaknesses.
IV-A A novel baseline, Lico-Net, for RPM problem
The RPM problem requires participants to select answers from an option pool in order to complete a $3\times 3$ progressive matrix of images. If the completed matrix adheres to a specific image progressive pattern that can be inferred from the incomplete matrix (problem statement), it is considered a correct answer selection[39]. Therefore, a successful RPM discriminator is expected to be able to identify the rationality of the progression patterns within a complete $3\times 3$ progressive matrix.
We require an outstanding baseline, leading us to propose the Lico-Net (link-conception net) as a novel network for RPM solvers. Lico-Net is designed to assess the rationality of the image progression patterns within a complete $3\times 3$ RPM matrix, and to represent this rationality through a score. Accordingly, we can sequentially employ all the options provided in an RPM case to complete the problem matrix, leverage Lico-Net to evaluate the progressive rationality of the matrix, and assign scores to the options based on this rationality. This approach allows us to reason and solve RPM problems effectively. The design details of Lico-Net are described as follows.
IV-A1 Perceptron module of Lico-Net
Aligned with RPM solver consensus on the importance of multi-viewpoint feature extraction[28, 23, 24, 32], Lico-Net uses the Vision Transformer (ViT) for feature extraction, keeping the entire output vectors, which are attention results, as multi-viewpoint features. We denote the number of viewpoints as L. Lico-Net then equally processes each viewpoint of extracted feature. The perceptron of Lico-Net is illustrated in figure 4.
Figure 4: Perceptron of Lico-Net
IV-A2 Reasoning module of Lico-Net
The Reasoning module of Lico-Net processes each view point feature generated by the perception module equally and in parallel, resulting in $L$ sets of inference outcomes. Subsequently, the mean value of these inference results is utilized as the global inference outcome for Lico-Net. The detailed process is described as follows.
For the equivalent viewpoint of an completed $3\times 3$ progressive matrix, we employ an MLP (Multi-layers perceptron) with a bottleneck structure to extract information from all minimal reasoning units (three images within a row for RAVEN and three images within a row or column for PGM) and obtain unit vectors. Specifically, we concatenate the representations of all images within a minimal reasoning unit in their inherent order and input them into the MLP. Compared to other structures, the MLP more directly preserves the order of the input representations and its structure is sufficiently simple. The extraction process of unit vector is illustrated in figure 5.
Figure 5: The extraction process of unit vector
Due to the design of RPM problems, typically multiple minimal reasoning units must be combined to arrive at the image progression patterns of a RPM matrix. For instance, a RAVEN case requires at least two row units to discern the case’s pattern, while PGM necessitates at least two row unites and two column unites[39]. It’s worth noting that determining the RPM matrix’s progression patterns don’t necessitate all minimal reasoning units; therefore, numerous combination methods of minimal reasoning units exist that are sufficient to identify the pattern. We refer to the combination requiring the least unit vectors as the minimal concept group.
Given that RPM problems might be constrained by multiple independent and decouple patterns, we embed $N$ optimizable vectors within each minimal concept group (in this paper, we embed two optimizable vectors, with $N$ being a hyperparameter that can be freely set) and input them into a standard Transformer-Encoder as tokens to compute $N$ kinds of potential progression patterns for the progressive matrix in the case. Assuming there are $M$ minimal concept groups, we have thus obtained $N\times M$ progression pattern vectors under one viewpoint. We denote the progressive pattern vector from a single perspective as $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$. These vectors have additional purposes to fulfill. This calculation process can be expressed in figure 6.
Figure 6: The extraction process of pattern vector
This paper independently calculates the scores of each option in the problem under $N$ distinct patterns.
The $M$ minimal concept groups can be divided into two categories: those involving options and those not involving options. In this paper, we select the pattern vectors calculated from the concept groups not involving options as queries, and the pattern vectors involving options as keys and values. We then compute the multi-head cross-attention results between them (using Transformer-Decoder) and use a new MLP to score the attention outputs. Since we intentionally introduced $N$ patterns, we obtain $N$ scores, which are summed to yield the final score for one viewpoint. This process is illustrated in figure 7.
Figure 7: The calculation process of final score for one viewpoint
Finally, we calculate the overall score of one option by averaging the scores from all viewpoints. Aligned with the answer logic, we employ the Cross-Entropy loss function to constrain these scores for optimizing Lico-Net. It’s worth noting that we treat each dimension of the unit vector as sample results from a distribution and use covariance matrices as additional loss function term to control the correlation between these distributions, aiming to encourage the unit vectors to contain more information. This approach bears resemblance to ViCReg[40].
In summary, after completing the RPM problem matrix with any given option, Lico-Net employs its perception module to encode each image within the problem matrix from multiple viewpoints, resulting in $L$ sets of $3\times 3$ image representations. The subsequent reasoning module of Lico-Net processes these $L$ sets of representations in parallel and equally. For any given set of representations, Lico-Net combines them to form all possible $M$ minimal concept groups. Subsequently, Lico-Net extracts $N$ progressive patterns from each minimal concept group, totaling $N\times M$ progressive pattern vector s. Finally, Lico-Net utilizes an cross-attention mechanism to measure the consistency of each progressive pattern expressed by the $M$ minimal concept groups and reflects this consistency through scores, resulting in $N$ consistency scores. The final score for the option is determined by averaging these $L\times N$ consistency scores.
IV-B The baseline for Bongard-Logo problem
The cases designers of Bongard-logo drew from high-dimensional human concepts and knowledge, carefully categorizing the Bongard-logo database into three distinct problem types: FF, BA, and HD. It is worth noting that the concept HD is decomposed into NV and CM during testing. This classification allows us to abstractly represent the distribution of the primary group in each Bongard-logo case as $p_{i}(x|\,y)$, while the distribution of the auxiliary group is denoted as $q_{i}(x|\,y)$. Here, $y$ denotes the reasoning type of the case, which can take values from the set FF, BA, HD, and $i$ represents the case index, ranging from $1$ to $n$, where $n$ is the total number of cases. Additionally, we denote the Bongard-Logo images as $x_{ij}$. Specifically, $\{x_{ij}|\,j\in[1,6]\}$ represents images in the $i$-th primary group, while $\{x_{ij}|\,j\in[8,13]\}$ represents images in the $i$-th auxiliary group. Additionally, $x_{i7}$ represents the test image to be potentially assigned to the $i$-th primary group, and $x_{i14}$ represents the test image to be potentially assigned to the $i$-th auxiliary group.
We copy a baseline from Triple-CFN [32], which refers to the baseline as $f_{\theta}(z|\,x)$ that maps samples $x_{ij}$ to latent variables $z_{ij}$. The baseline employs the InfoNCE loss function, which results in the $f_{\theta}(z|\,x)$ being able to solve Bongard-logo to a certain degree. This approach fosters a higher degree of similarity between $z_{ij}$ and $z_{i\tilde{j}}$ (where $\tilde{j}\neq j$ and $\tilde{j}$, $j$ $\in[1,7]$), compared to the similarity between $z_{ij}$ and $\{z_{ij}|\,j\in[8,13]\}$, thereby emphasizing the distinction between primary and auxiliary representations. This process aligns with the underlying logic of Bongard-logo.
This paper reproduces the baseline experiments in Triple-CFN [32], which were originally conducted on the Bongard-Logo using the ResNet18 model. The baseline focuses on learning four concepts in Bongard-Logo — ‘FF’, ‘BA’, ‘NV’, and ‘CM’ — separately or in combination [32]. To assess the impact of model size on the baseline experimental findings, we have replaced the backbone of the baseline experiments from ResNet18 to ResNet50 and repeated the experiments. The experimental results are presented in Table I.
The findings indicate that when presented with a combination of databases, both ResNet18 and ResNet50 experience confusion among the four distinct human concepts. The varying degrees of decrease in the models’ reasoning accuracy on the FF and BA problems, accompanied by an increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion. Furthermore, this confusion cannot be attributed to insufficient network capacity, as even ResNet50, which is more capacious than the ResNet18 used in Triple-CFN’s baseline, is subject to the same challenge.
TABLE I: Reasoning Accuracies of DC based ResNet based ResNet50 on Bongard-logo.
Test Accuracy(%)
Model and Data Set FF BA CM NV
ResNet18 on Bongard-logo 88.1 97.9 76.0 75.8
ResNet18 on Separated Bongard-logo 97.9 99.0 75.0 72.8
ResNet50 on Bongard-logo 88.9 98.0 76.1 76.0
ResNet50 on Separated Bongard-logo 98.3 99.2 75.3 72.2
After conducting backbone substitution experiments, this paper aligns with the perspective of Triple-CFN [32]. Triple-CFN argues that the identification of image attributes in the Bongard-Logo problem relies heavily on primary and auxiliary samples in specific cases [32], posing challenges due to this context-dependent attribute determination. In other words, the reasoning identity of these images is contingent upon context-specific concepts, leading to multifaceted interpretations of image attributes. Consequently, deep networks encounter confusion when processing high-dimensional Bongard-Logo concepts [32]. In contrast, the RPM problem exhibits consistency in its recognition of image attributes, such that a white medium-sized triangle retains its identity regardless of its appearance in any RPM case option pool.
V Distribution Distance-based Discriminator of Case (D3C)
Using distributions to describe human concepts in abstract reasoning problems for deep neural networks is advantageous compared to the use of vectors. This is because distributions present a more comprehensive and nuanced means of capturing the inherent diversity and uncertainty in human concepts. Human concepts are frequently characterized by ambiguity, polysemy, intersectionality, and overlap, making their representation via static vectors challenging. Conversely, distributions afford a dynamic and probabilistic framework that can effectively accommodate such inherent variability. Employing distributions to represent human concepts endows deep neural networks with richer and more flexible information. This, in turn, enhances their capability to comprehend and resolve complex, uncertain abstract reasoning problems. Probability distributions illustrate the possibilities and variations associated with a concept, whereas Gaussian distributions specifically model the underlying structure of a concept.
V-A D3C for Bongard-Logo
D3C argues that distributions are better abstract concepts than vectors. Therefore, it employs a deep network $f(z|\,x)$ to estimate concepts in Bongard-Logo images and represents them as distributions. By measuring concept distances as the distances between these distributions, the method becomes applicable to solving the Bongard-Logo problem. D3C is a plug-and-play method for any network with form of $f(z|\,x)$, improving models like Triple-CFN and ResNet50.
In practice, this paper assumes the concepts in Bongard-Logo cases follow a Gaussian distribution. We use a backbone network $f(z|\,x)$ to extract these concepts as distributions. To do this, we adapt the network to output $\mu$ and $\sigma$ instead of a single vector $z$, turning $f(z|\,x)$ into $f(\mu,\sigma|\,x)$. The vector $z$ is split into two equal-length segments $\mu$ and $\log(\sigma^{2})$, representing mean and variance logarithm, respectively. This method normalizes original outputs $z$ to Gaussian distributions $\{\mu,\sigma^{2}\}$ and bears resemblance to the Variational Autoencoder (VAE) framework[8]. Representing concepts as distributions, not vectors, provides more flexibility and accuracy in abstract reasoning.
With the estimated concepts for each image in the Bongard-Logo dataset represented as distributions, we compute the Sinkhorn distance between these distributions using the reparameterization technique[8]. Subsequently, we impose constraints on these distances to train and optimize the backbone network $f(z|\,x)$, aligning with the underlying logic of the cases. The optimization process involves minimizing a binary cross-entropy loss, which is calculated using the exponential of the negative Sinkhorn distance as a factor. The entirety of this process is illustrated in Figure 8. When applied to a batch of Bongard-Logo cases, the D3C method can be summarized as follows:
${\ell_{\text{D3C}}}(\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\})$
$=\sum_{i=1}^{b}\sum_{j=1}^{7}\sum_{\tilde{j}=j+1}^{7}{\text{S}(\{z_{ijs}|\,s\in[1,S]\},\,\{z_{i\tilde{j}s}|\,s\in[1,S]\})}$
$+\sum_{i=1}^{b}\sum_{{j}=1,}^{7}\sum_{\tilde{j}=7}^{14}\log({1-{e^{-\text{S}(\{z_{ijs}|\,s\in[1,S]\},\,\{z_{i\tilde{j}s}|\,s\in[1,S]\})}}})$ (2)
[…]
Figure 8: Feedforward process of D3C
The design of D3C allows it to be applied to any image encoder that can be abstracted as $f(z|x)$. A wide range of encoders, including those from the ResNet family, ViT, and specialized encoders like Triple-CFN—which is particularly suitable for Bongard-Logo problems—meet the abstraction requirement and are thus compatible with D3C.
V-B D3C on Lico-Net for RPM
Notably, in contrast to the Bongard-Logo database, each case in RAVEN and PGM is accompanied by well-defined and comprehensive rule labels in the form of Meta data. These auxiliary labels provide descriptions of the image progression patterns within the cases, facilitating the delineation of concept boundaries. Therefore, this paper decides to introduce the Meta data from the RPM database into the Lico-Net, aiming to organize the distribution of the encoded progressive pattern vector $\{P_{nm}\}$ in Lico-Net. This enhancement is expected to improve both the reasoning accuracy and human-perspective interpretability of Lico-Net. This process will be implemented by the D3C method.
Within the frameworks of RAVEN and PGM, there exists a many-to-one mapping between image progression patterns and their corresponding concepts. For instance, both single-step and double-step increments, as well as single-step and double-step decrements in image attributes, are categorized under the umbrella of the “progressive” concept according to Meta data. Therefore, this paper proposes extracting the concepts described by the Meta data and representing them as distributions, which aligns with the original idea behind D3C. D3C posits that using distributions to describe concepts is more suitable than using vectors. As Lico-Net has already extracted the progression patterns of images for the RPM matrix and encoded them as vectors, we can naturally employ concepts with distribution forms to constrain and delimit these image progression pattern vectors, ensuring that these progression pattern vectors belong to the distribution of corresponding concept. This approach will optimize Lico-Net’s learning of rules and clarify the boundaries among various concepts. The aforementioned idea can be represented in the framework illustrated in the Figure 9.
Figure 9: The framework of D3C
In practice for D3C, we utilized a standard Transformer-Encoder to convert the textual Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data. Specifically, the initial portion of the concept vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the concepts described in Meta data and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$. Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ termed as progression pattern vectors. Lico-Net extracts N types of progression patterns (where N can be freely set) for each complete RPM matrix, and encodes M progression pattern vectors (where M is determined by the number of existing minimal concept groups) for each progression pattern, forming the vector group $\{P_{nm}\}$. This process can be reviewed in Figure 6. To integrate Lico-Net with the D3C framework, the number of encoding patterns $N$ in Lico-Net, which can no longer be set arbitrarily, is set according to the instance of the RPM problem, depending on the number of decoupled concepts present in the Meta data of that instance. For example, in PGM, there are two such concepts: shape pattern and line pattern, and $N$ is set to 2. The special setting of $N$ in Lico-Net can be viewed as $N$ slots, which can be aligned with the $N$ decoupled concepts described in the Meta data, respectively. Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns $\{P_{nm}\}$ should exhibit a high likelihood of adhering to the distribution outlined by their corresponding concepts, and conversely, demonstrate a low probability of deviation.
To impose constraints on the distributions of the extracted $N$ progression patterns, we have devised the following process. By synthesizing all perspectives—specifically, by calculating the mean of the progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ across all $L$ viewpoints—Lico-Net encodes $N\times M$ vectors, denoted as $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$, representing the full-viewpoint progression patterns. The set $\{\overline{P}_{nm}\}$ can be interpreted as the outcome of sampling $M$ times from each of the $N$ distributions of progression patterns across all viewpoints generated by Lico-Net. This interpreting perspective enables us to utilize Meta data to enforce constraints on the $N$ distributions. Next, we sample $S$ times from the distribution of concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ to obtain sampling results $\{r_{ks}|\,k\in[1,K],s\in[1,S]\}$. Consequently, this paper employs the Sinkhorn distance to constrain the distance between the distributions of the $N$ kinds of progression patterns encoded by Lico-Net and their corresponding concept distributions in the Meta data. The process of aligning $\{\overline{P}_{nm}\}$ with the distribution of concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ can be illustrated in figure 10.
Figure 10: Feedforward process of D3C
The calculate porcess can be exprecessed as:
${\ell_{\text{D3C}}}(\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\},\{r_{ks}|\,k\in[1,K],s\in[1,S]\})$
$=\sum_{n=1,\tilde{k}|\,n}^{N}{\text{S}(\{\overline{P}_{nm}|\,m\in[1,M]\},\,\{r_{\tilde{k}s}|\,s\in[1,S]\})}$
$+\sum_{k=1,k\neq\tilde{k},\tilde{k}|\,n}^{K}\sum_{n=1}^{N}\log({1-{e^{-\text{S}((\{\overline{P}_{nm}|\,m\in[1,M]\},\,\{r_{{k}s}|\,s\in[1,S]\}))}}})$ (3)
[…]
In summary, Lico-Net is capable of extracting $N$ types of progression patterns for each minimal concept group in the RPM problem. Given the existence of $M$ minimal concept groups, we obtain a total of $N\times M$ progression pattern vectors. To constrain and clarify the distribution of vectors under each of the $N$ progression patterns, we consider the vectors as the outcome of $M$ samplings from each of the $N$ distributions. This perspective enables us to implement distributional constraints based on concepts in Meta data, thereby enhancing the alignment and interpretability of the progression patterns extracted by Lico-Net.
V-C A low-cost loss function for the D3C, the Hermes for the Zeus
Sampling $S$ times (where $S$ is arbitrarily set, and $S$ is set to $20$ for Sinkhorn distance) from the constructed distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of concepts in Meta data and yielding $\{r_{ks}|\,k\in[1,K],s\in[1,S]\}$, we calculate the Sinkhorn distance between the progression pattern vectors and these sampling results $\{r_{ks}\}$. Typically, this approach requires a substantial number of samples and incurs a considerable computational cost. To reduce computational cost and sampling frequency, this paper employs the mean of cosine similarity between sampling vectors as an alternative distance metric norm to the Sinkhorn distance.
The calculate process of new method applied on Bongard-Logo can be expressed as:
${\ell_{\text{D3C-cos}}}(\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\})$
$=-\sum_{i=1}^{b}\sum_{j=1}^{7}\sum_{\tilde{j}=j+1}^{7}\sum_{s=1}^{S}\log\frac{{{e^{({z_{ijs}}\cdot{z_{i\tilde{j}s}})/t}}}}{{{e^{({z_{ijs}}\cdot{z_{i\tilde{j}s}})/t}}+\sum\nolimits_{k=8}^{14}{{e^{({z_{ijs}}\cdot{z_{i{k}s}})/t}}}}}$ (4)
The calculate process of new method aligned with Lico-Net can be expressed as:
${\ell_{\text{D3C-cos}}}(\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\},\{r_{ks}|\,k\in[1,K],s\in[1,S]\})$
$=-\sum_{n=1,\tilde{k}|\,n}^{N}\sum_{m=1}^{M}\sum_{s=1}^{S}\log\frac{{{e^{({\overline{P}_{nm}}\cdot{r_{\tilde{k}s}})/t}}}}{{{e^{({\overline{P}_{nm}}\cdot{r_{\tilde{k}s}})/t}}+\sum\nolimits_{k=1,k\neq\tilde{k}}^{K}{{e^{({\overline{P}_{nm}}\cdot{r_{{k}s}})/t}}}}}$ (5)
The $t$ is set to $10^{-6}$ and $S$ is set to $5$ in function (V-C) and (V-C). The logic of symbols in function (V-C) and (V-C) has mentioned. It is evident that, compared to function (V-A) and (V-B), the demand for sampling frequency $S$ in D3C-cos has been significantly reduced.
Estimating distribution distance via cosine similarity between samples can be seen as a modified Wasserstein distance, relinquishing lower bound estimation and having a fixed Lipschitz constant. This ”D3C-cos” method, while not exact, provides a cost-effective loss function for distribution distance, acting as a substitute for the Sinkhorn distance. We still use Sinkhorn distance for performance testing. The success of such simplification can be attributed, in part, to our modeling of the distribution of pattern descriptions as a Gaussian distribution.
VI Dynamic Delineation of the Concept (D2C)
In this section, the paper initially endeavors to introduce novel constraints for the Bongard-Logo discriminators, aiming to enhance their performance. Motivated by these constraints, we incorporate innovative designs into the proposed Lico-Net architecture.
VI-A D2C on Bongard-Logo problem
In the Bongard-Logo problem, the determination of image attributes is influenced by the specific contextual concepts associated with each individual case [32]. Distinct concepts require the network to encode specific pixel configurations into different attribute representations, potentially compelling the network to maintain mutually exclusive representation encoding strategies and methods [32]. This scenario often leads to concept confusion within the network, rendering it uncertain about which concept to rely on when processing a Bongard-Logo case [32]. In a nutshell, from the perspective of network, there exists mutual infringement and erroneous coupling among the high-dimensional human concepts in the Bongard-Logo cases. Therefore, this study aims to develop a methodology that effectively decouples the concepts across different cases within the Bongard-Logo problem. We posit that by successfully decoupling such that distinct cases rely on distinct concepts we may alleviate, to a certain extent, the issue of the network losing concept matching when confronted with a Bongard-Logo problem.
Therefore, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases. By establishing a one-to-one correspondence between problem cases and concepts, we aim to enable the network to avoid uncertainty in determining which concept to rely on when solving a specific case. This approach aims to mitigate the impact of concept confusion on the solution for the Bongard-Logo problem.
To foster concept antagonism between different Bongard-Logo cases, we have revised the data sampling techniques employed in mini-batch training methodologies. Furthermore, an additional loss function term has been devised to facilitate the reassignment of concepts within the network architecture. This approach, termed “D2C” (Dynamic Delineation of Concepts), endeavors to cultivate a state of concept contrast amidst diverse Bongard-Logo cases.
VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
The design idea of D2C phase 1 is quite simple, which involves introducing primary group samples from other cases as additional samples into the auxiliary group of the current case. During the implementation process, for each case under consideration, we introduce two distinct cases into it. This process can be regarded as ancestral sampling. This constitutes a plug-and-play approach for training most Bongard-Logo solver such as PMoC[31] and Triple-CFN[32], which is designated as phase 1 of the D2C method in this paper. The process of D2C phase 1 is in Figure 11. As evident from the figure, the ancestral sampling process can be accomplished by performing a misalignment stacking of the sampled Bongard-Logo cases. By incorporating this additional ancestral sampling procedure into the traditional mini-batch training sampling methods, this study constructs a Bongard-logo problem with reorganized concept space for training purposes. This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts. This paper aims to utilize this strategy to mitigate concept infringement and erroneous coupling among different Bongard-Logo cases.
Figure 11: Feedforward process of D2C phase 1
It is worth emphasizing that the improvement in model performance provided by D2C is not solely attributed to an increase in the number of negative examples within each contrastive group. This paper posits that the incorporation of negative examples, which assist in delineating concept boundaries among cases, holds the key to enhancing a model’s proficiency in abstract reasoning tasks. Simply adding unhelpful or irrelevant negative examples to the dataset does not significantly improve model performance. In an attempt to enrich the dataset, this study generated a substantial amount of Bongard-logo image data using DDPM [9] and integrated these as auxiliary group sample within the original problems. However, this methodology based on DDPM [9] did not significantly enhance model accuracy.
VI-A2 D2C phase 2: The concept delineation method based on a new loss function term
This paper proposes that, besides changing training data sampling, adding a loss function to limit the similarity of primary group samples’ representations across cases in a Bongard-Logo training batch implicitly fulfills the D2C’s main target.
Specifically, in this paper, we exclusively retain the representation of primary group within the set $\{z_{ij}|\,i\in[1,b],j\in[1,14]\}$, which is $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$. Subsequently, we randomly partition the primary representation of each case within $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ into two distinct subsets: $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$. By taking the mean of the vectors within $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, we derive the corresponding vectors $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$. Ultimately, we compute InfoNCE loss function for constraining the similarity among $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$ as the additional loss function item. The specific calculation process can be explained by:
${\ell_{\text{D2C}}}(\{z_{ij}|\,\,i\in[1,b],j\in[1,7]\})$
$=-\sum^{b}_{i=1}\log\frac{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}}}{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}+\sum\nolimits_{\tilde{i}=1,\tilde{i}\neq i}^{b}{{e^{({v_{i}}\cdot{v_{\tilde{i}}})/t}}}}}$ (6)
$\text{where}\,\,v_{i}=\sum\{z_{ih}\},\,\tilde{v}_{i}=\sum\{z_{i\tilde{h}}\}\,\,$
$\text{and}\,\,\{z_{ih}\}\cup\{z_{i\tilde{h}}\}=\{z_{ij}|\,j\in[1,7]\}$
This loss term is to decrease the similarity of primary group representations across different cases. Evidently, D2C employs a ”rather be safe than sorry” approach, which may lead to some overcorrection.
Essentially, D2C’s phase 2 aims to achieve a similar goal as the phase 1 by introducing a new loss function. However, this makes phase 2 mainly compatible with backbones abstractable as $f(z|x)$. This allows most image encoders like ResNet and Triple-CFN [32] to be compatible with full D2C. But PMoC[31], relying on probability measures, is limited to D2C’s phase 1.
VI-B D2C on RPM probelm
Lico-Net can achieve the integration with D2C by introducing images from other RPM cases as incorrect options within the option pool of the current RPM case. It’s worth noting that the application of D2C in RPM discriminators is confined to phase 1 due to the complexity arising in the implementation of phase 2. This process is shown in Figure 12.
Figure 12: Feedforward process of Lico-Net aligned with D2C
VI-C The summary of D2C
It is evident that D2C employs a precautionary approach, aiming to minimize mutual infringement and erroneous coupling among concepts, even at the cost of potential overcorrection. However, we posit that D2C is suitable for addressing the Bongard-Logo problem, where there exists mutual infringement and erroneous coupling of concepts across cases, but it is challenging to apply D2C to the RPM problem, which involves sharing a common set of concepts across cases and focuses on consistent attribute identification.
VII Defense-Driven Discrimination of Distributed Concepts (D4C)
The performances of D2C and D3C on both Bongard-Logo and RPM tasks have stimulated deeper reflections on abstract reasoning challenges. This raises some fascinating questions: Why does modifying negative examples have a profound impact on the solver’s performance specifically in the Bongard-Logo problem? Beyond posing difficulties for the discriminator, what are the additional roles played by positive and negative examples in abstract reasoning challenges? After thoughtful consideration of these aspects, this section introduces the D4C approach.
VII-A Negative samples are not merely intended to confound the discriminator
This paper posits that it is worthwhile to consider the significance of positive and negative examples in abstract reasoning problems. Currently, existing discriminative models for solving RPM and Bongard-Logo problems often adhere to a pattern known as the probability labeling scheme. As the name suggests, when a positive example is submitted to the reasoning problem’s statement, the discriminator will respond with a high probability, whereas the negative examples will be labeled with a low one. We can view this pattern in the following way: the discriminator constructs a solution distribution for the cases presented in abstract reasoning problems. Within this distribution, positive examples are assigned higher probabilities, whereas negative examples are assigned lower ones. The endeavor to train the discriminator can be conceptualized as a process of optimizing and fine-tuning the solution distribution, with the ultimate objective of aligning it more closely with the underlying distribution of correct solutions for the given cases.
To further explain our approach in a clearer and intuitive manner, we assume that the distribution of correct solutions for each case in a reasoning problem follow a Gaussian distribution. Given that some reasoning problems may have multiple correct solutions, it is also possible that the distribution is a mixed Gaussian distribution. Using positive and negative examples to estimate the parameters of a Gaussian distribution through the cross-entropy loss function can be seen as leveraging them to optimize the following function:
$-\log(p(x,\mu,\sigma))=\frac{{(x_{i}-\mu_{i})^{2}}}{{2\sigma^{2}_{i}}}-\log(|\,\sigma_{i}|\,)-\frac{1}{2}\log(2\pi)$ (7)
Where $x$ represents the positive or negative samples, $\mu$ and $\sigma$ represent the mean and standard deviation of the Gaussian distribution separately. It can be observed that when positive samples are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the Euclidean distance between sample $x$ and the mean of the correct solution $\mu$. Conversely, when negative samples are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$.
Therefore, this paper posits that in reasoning problems, the establishment of positive examples aids deep models in estimating the mean of the correct solution distribution, while negative examples serve to limit the variance. The variance represents the distance between the boundaries of the distribution and its center. Correspondingly, we posit that both positive and negative examples play a crucial role in abstract reasoning tasks. Positive examples assist in determining the centroid of the concepts presented in each case, whereas negative examples contribute to delineating the boundaries of these concepts. Thus, whether negative examples can minimize the variance $\sigma^{2}$ to the size of the correct distribution becomes crucial. If the quality of negative examples is not high enough, the variance cannot be minimized effectively, in other words, it cannot delineate a more precise concept boundary. This underscores the significance of establishing both positive and negative examples in abstract reasoning problems.
This significance reflects a despairing fact that the upper limit of the discriminator’s reasoning accuracy in this abstract reasoning problem is determined by the quality of positive and negative examples in the question.
VII-B High-quality negative samples lead to more rigorous concept boundaries
The aforementioned perspective is the core contribution of this paper. Since negative examples delineate the boundaries of concepts, we posit that enhancing the quality of negative examples equates to delineating more precise concept boundaries. A more precise boundary imply superior discriminator performance. Therefore, this paper aims to device a method that can enhance the quality of negative examples in reasoning problems, and D2C is one such approach.
From the novel perspective on negative samples in reasoning problems, the D2C methodology presented in this study essentially defines a new problem boundary for deep networks. However, we suggest that this approach of defining a new concept boundary, specifically the introduction of images from other cases as negative examples for the current case, can make a significant contribution to problems like Bongard-Logo, where there exists concept infringement and erroneous coupling among different cases. Conversely, for problems like RPM, where different cases share a common concept set, this method may not contribute effectively. In other words, due to the specific characteristics of the Bongard-Logo problem, D2C’s approach has, to some extent, enhanced the quality of negative examples and delineated more precise concept boundaries. However, it cannot replicate these improvements in RPM cases, owing to the distinct problem characteristics of RPM.
In a nutshell, while D2C introduces new negative examples and concept boundaries for reasoning problems, it cannot guarantee the superior quality of these examples or the precision of the new concept boundaries. Hence, this paper proposes the D4C (Defense-Driven Discrimination of Distributed Concepts) method, which deliberately seeks out higher-quality negative examples for reasoning problems, thereby defining more accurate concept boundaries. The underlying idea of D4C is to generate higher-quality negative samples for abstract reasoning problems, aiming to delineate a more appropriate problem boundary for the discriminator compared to the original problem setup in abstract reasoning. By leveraging novel mechanisms and strategies, D4C strives to deliver more precise and reliable definitions of concept boundaries, fostering further advancements in the realm of abstract reasoning.
VII-C Impelmention of D4C method
The D4C methodology proposed in this paper emphasizes on devising a more suitable concept boundary for both RPM and Bongard-Logo discriminators, aiming to enhance their problem-solving capabilities. To achieve this objective, a distinctive approach centered on the generation of confusing samples is introduced. These samples are intentionally designed to challenge and improve the discriminator’s accuracy. A dedicated generator has been devised to carry out this generation process. Through iterative defense against attacks launched by this generator, the discriminator progressively refines its discrimination and classification skills, leading to overall performance enhancement in both areas. Specifically, within this study, a generator has been tailored to augment the capabilities of two distinct Bongard-Logo discriminators: PMoC and Triple-CFN. Additionally, this generator can also be directly utilized for the refinement of Lico-Net. Nevertheless, acknowledging the inherent disparities between RPM problems and Bongard-Logo challenges, additional designs have been incorporated into the generator to cater more effectively to the refinement requirements of Lico-Net.
In this section, we delve into the design of the discriminator and generator within the D4C framework as it pertains to the Bongard-Logo and RPM problems.
VII-C1 discriminator for Bongard-logo
Effective Bongard-Logo discriminators abound, with our preceding works, such as PMoC[31] and Triple-CFN[32], achieving notable successes. Regarding discriminators capable of handling adversarial tasks, the discriminator must undertake the task of measuring the distance between samples generated by the generator and real samples. This necessitates that the discriminator satisfies Lipschitz continuity to a certain extent[36]. Therefore, in this paper, we choose the spectrum-normalized PMoC and Triple-CFN as discriminators for our D4C, and design generators to refine them.
In this study, the choice of PMoC, constrained by spectral normalization, as the discriminator for the D4C method on Bongard-logo is motivated by its explicit measurement of the distance between auxiliary and primary group sample distributions. Conversely, the task of Triple-CFN is relatively more intricate as it considers not only the distance between primary and auxiliary samples but also the distance among primary samples. Consequently, compared to PMoC, this study entails more elaborate designs for the generator of Triple-CFN.
VII-C2 generator for Bongard-logo
In this paper, we argue that instead of forcing the generator to produce samples that can confuse the discriminator, it is more effective to generate representations of confusing samples, referred to as confusing representations. This approach reduces the computational burden on the network, as measured by intuitive metrics such as computational complexity.
For the Bongard-Logo problem, we model the overall distribution as a mixture of Gaussian distributions. Within this framework, each individual primary group in a Bongard-Logo case is represented as a separate Gaussian distribution. Consequently, the D4C generator designed for the Bongard-Logo problem exhibits a relatively straightforward structure. The generator aims to encode a distribution based on the representations of the primary groups in a given Bongard-Logo case, ensuring that any representation sampled from this distribution is capable of effectively confusing the discriminator.
Specifically, we employ a standard Transformer encoder to encode a Gaussian distribution based on the representations of the primary samples in a Bongard-Logo case, along with an optimizable vector. This encoding approach bears similarity to D3C, wherein we extract the optimizable vector from the attention results and process it through a linear feed-forward layer to obtain the mean and logarithm of variance for a Gaussian distribution. Subsequently, we sample multiple representations from this distribution as confusing representations to challenge the discriminator, which must strive to counter them. The generator’s training objective is to generate representations for Bongard-Logo cases that achieve higher scores in the discriminator compared to the representations of the original negative examples in the case. Conversely, the discriminator’s training objective is unambiguous: to classify these generated representations as negative examples. The aforementioned approach can be seamlessly applied to the PMoC model, providing a straightforward and effective means of enhancing its performance. The integration method between the generator and PMoC is illustrated in Figure 13.
Figure 13: Feedforward process of D4C aligned with PMoC on Bongard-Logo
For the generator in Triple-CFN[32], it is essential to decay the generated representations. Specifically, when the generator with the aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation must be weighted and summed with the representations in the auxiliary group. After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to decay the confusing representation. Due to the mechanism of Triple-CFN, this decay process is deemed necessary, which defines an image’s reasoning identity based on the direction of its representation. This mechanism makes it easier for a Transformer Encoder-based generator to generate realistic confusing representations, potentially causing Triple-CFN to collapse. Therefore, we have designed the aforementioned decay process. The integration method between the generator and Triple-CFN is illustrated in Figure 14.
Figure 14: Feedforward process of D4C aligned with Triple-CFN on Bongard-Logo
VII-C3 discriminator of RPM
In addressing Raven’s Progressive Matrices (RPM) problems, this study places significant confidence in the Lico-Net model, considering it as an exceptional benchmark. It is anticipated that the application of spectral normalization will enhance the performance of Lico-Net, positioning it as a highly suitable discriminator for the D4C methodology.
Lico-Net will encounter challenges posed by the generator, which produces confusing representations aimed at completing the $3\times 3$ reasoning matrix. To address these challenges, Lico-Net will undertake an assessment of the internal consistency within the progressive patterns of the $3\times 3$ reasoning matrix, which are completed by these confusing representations. Any inconsistencies or irregularities identified within these representations will be flagged as negative samples, indicating incorrect responses. Through this rigorous process, Lico-Net strives to accurately identify and reject confusing representations, thereby augmenting its overall precision and reliability.
VII-C4 generator of RPM
Unlike the Bongard-Logo problem, the answer to a single case in the RPM problem exhibits non-uniqueness, and thus can be represented as a mixture of Gaussian distributions. Therefore, our aspiration is to tailor a generator for the D4C method within the RPM context, capable of encoding a distribution that reflects this mixed Gaussian nature. In this paper, we employ a glove-shaped network to simulate the process of sampling from such mixed Gaussian distributions.
Specifically, we utilize a standard Transformer-Encoder to process the representations of all images in the statement of an RPM case and $N$ pattern vectors from the statement encoded by Lico-Net, alongside $W$ optimizable vectors (where $W$ is a hyperparameter) and a stochastic promoter. The promoter is a vector sampled from a multivariate Gaussian distribution, wherein the mean is optimizable while the variance is fixed at 1. The Transformer-Encoder needs to process $N+W+1+8$ vectors in total. Notably, when these $W$ vectors and the promoter are arranged together, they resemble the shape of a glove.
It is worth noting that we have devised a novel positional encoding embedding approach for the Transformer-Encoder, which serves as the backbone of our RPM generator. This encoding aims to restore the $3\times 3$ image arrangement characteristic of the RPM problem matrix. For the PGM problem, our positional encoding scheme assigns unique encodings to images located on the diagonal and upper triangular regions within the problem matrix of a given PGM case. Correspondingly, the positional encodings for images in the lower triangular region are symmetrical with respect to the diagonal, mirroring the encodings of their counterparts in the upper triangular region. However, it should be acknowledged that this paper does not demonstrate the superiority of D4C on the relatively less challenging RAVEN dataset.
Moving on to the main topic, after obtaining attention results from the Transformer-Encoder, we extract $W$ optimizable vectors from these results and process them into $W$ distributions using a linear layer, like D3C method. Following the extraction of the promoter, it is processed into a probability output using a linear layer and hard Gumbel-Softmax activation, yielding highlighted probabilities for the $W$ distributions. Subsequently, we sample representations, referred to as confusing representations, from the highlighted distributions each time. In the technical implementation, we employ the reparameterization trick to draw samples from each distribution, resulting in $W$ sampling outcomes. We then perform a weighted sum of these sampling results using the hard-highlighted probabilities processed by the hard Gumbel-Softmax, thus completing the entire mixed-Gaussian sampling process. The methodology outlined in this section is coined as “D4C glove-train”. The forward process of D4C on Lico-Net is shown in the figure 15. In figure 15, “pattern extractor” is illustrated in figure 6 and the process of “pattern scoring” is shown in figure 7.
The training objective of the glove-shaped generator remains consistent: to produce confusing representations that achieve higher scores in the discriminator compared to the original negative examples provided by the dataset. Conversely, the discriminator aims to resist the interference posed by these confusing representations.
Figure 15: Feedforward process of D4C on RPM problem
Overall, compared to traditional methods, D4C is a fundamentally different approach. Previous studies on solving abstract reasoning problems[23, 21, 28, 32] have primarily focused on constructing solvers that structurally align with the inductive bias in the problems, including our Lico-Net. However, our D4C reconsiders the significance of negative examples in reasoning problems, and designs a solution based on this introspection. As a bonus, the glove-shaped generator in D4C provides suggestions for simulating the random sampling process of a mixed Gaussian distribution.
VIII Experiment
All experiments conducted in this paper were programmed in Python, utilizing the PyTorch[43] framework.
VIII-A Experiment on RAVEN
Initially, we tested Lico-Net on RAVEN and I-RAVEN datasets to show its potential. For a fair comparison, we used the same settings and equipment as RS-Tran[28] and Triple-CFN[32], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
TABLE II: Reasoning Accuracies on RAVEN and I-RAVEN.
Test Accuracy(%)
Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
SAVIR-T [24] 94.0/98.1 97.8/99.5 94.7/98.1 83.8/93.8 97.8/99.6 98.2/99.1 97.6/99.5 88.0/97.2
SCL [23, 24] 91.6/95.0 98.1/99.0 91.0/96.2 82.5/89.5 96.8/97.9 96.5/97.1 96.0/97.6 80.1/87.7
MRNet [21] 96.6/- -/- -/- -/- -/- -/- -/- -/-
RS-TRAN[28] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
Triple-CFN[32] 98.9/99.1 100.0/100.0 99.7/99.8 96.2/97.5 99.8/99.9 99.8/99.9 99.9/99.9 97.0/97.3
Triple-CFN+Re-space[32] 99.4/99.6 100.0/100.0 99.7/99.8 98.0/99.1 99.9/100.0 99.9/100.0 99.9/99.9 98.5/99.0
Lico-Net 99.5/99.7 100.0/100.0 99.7/99.8 98.2/99.2 99.9/100.0 99.9/100.0 99.9/99.9 98.8/99.2
An interesting observation from the Table II indicates a commonality among existing excellent RPM solvers, including Lico-Net. Specifically, these solvers exhibit varying reasoning accuracy when tackling the RAVEN [15] and I-RAVEN [37] datasets. The I-RAVEN dataset was designed to address certain limitations in the option pool configuration of RAVEN by introducing improvements to its negative example setup. These enhancements include, but are not limited to, increasing the number of attributes that are altered in the negative examples and augmenting the magnitude of attribute deviations within them. However, from a human perspective, since the setup of positive examples remains unchanged, such additional setups do not alter the difficulty of reasoning, making I-RAVEN neither more nor less challenging than RAVEN. This paper speculates that the disparate outcomes achieved by these models on RAVEN and I-RAVEN could potentially be attributed to the higher quality of negative examples present in the I-RAVEN dataset compared to those in RAVEN. Such phenomena echo the D4C’s spirit.
VIII-B Experiment on Bongard-Logo
We conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN [32] as backbone for D2C, D3C(-cos), and D4C and using PMoC [31] as backbone for D2C and D4C. PMoC, a probabilistic labeling model, does not meet the requirements of D3C(-cos), as D3C(-cos) demands a backbone in the form of $f(z|x)$. The experimental results are presented in the table III.
Observing table III, D2C and D3C(-cos) notably contributed to the Bongard-Logo task.Replacing D3C with D3C-cos when integrated with ResNet50 has minimal impact on performance. From the result-oriented perspective, D3C-cos is a qualified alternative for D3C. We didn’t test Triple-CFN with D3C or D2C+D3C due to practical limitations from our equipment, which led to the proposal of D3C-cos. Furthermore, D4C has significantly improved Triple-CFN, ResNet50, and PMoC, often matching the combined effects of D2C and D3C. However, on Bongard-Logo’s FF concept, D4C’s improvements on Triple-CFN and ResNet50 were less than ideal. This might be because D4C was designed for probabilistic labeling models, whereas Triple-CFN and ResNet50 do not follow this probabilistic approach. A decay mechanism, as a compromise for non-probabilistic labeling models, was added when integrating D4C with these models. Experimental results suggest this mechanism needs further refinement. Future work should explore better designs to adapt D4C to non-probabilistic labeling models like Triple-CFN. Conversely, PMoC [31], a probabilistic labeling model for Bongard-Logo, benefits more from D4C due to its scheme alignment.
TABLE III: Reasoning Accuracies of models on Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
SNAIL 59.2 56.3 60.2 60.1 61.3
ProtoNet 73.3 64.6 72.4 62.4 65.4
MetaOptNet 75.9 60.3 71.6 65.9 67.5
ANIL 69.7 56.6 59.0 59.6 61.0
Meta-Baseline-SC 75.4 66.3 73.3 63.5 63.9
Meta-Baseline-MoCo[42] 81.2 65.9 72.2 63.9 64.7
WReN-Bongard 78.7 50.1 50.9 53.8 54.3
SBSD 83.7 75.2 91.5 71.0 74.1
PMoC 92.0 90.6 97.7 77.3 76.0
Triple-CFN 93.2 92.0 98.2 78.0 78.1
ResNet50 (infoNCE baseline) 92.0 88.9 98.0 76.1 76.0
ResNet50+D2C 92.6 90.5 98.2 77.0 77.2
ResNet50+D3C 92.3 90.6 98.3 76.5 76.8
ResNet50+D2C+D3C 93.0 91.0 98.5 77.5 77.5
ResNet50+D3C-cos 92.1 90.2 98.2 76.9 76.4
ResNet50+D2C+D3C-cos 92.9 90.8 98.4 77.5 77.4
ResNet50+D4C 93.2 90.8 98.5 77.2 77.3
Triple-CFN+D2C 93.9 92.7 98.5 78.2 78.5
Triple-CFN+D3C-cos 93.7 92.5 98.3 78.2 78.2
Triple-CFN+D2C+D3C-cos 94.1 93.0 98.5 78.2 78.6
Triple-CFN+D4C 94.3 92.8 98.8 78.7 79.0
PMoC+D2C 92.5 91.2 98.0 77.7 76.7
PMoC+D4C 93.0 92.6 98.2 78.0 77.5
Finally, this paper notes that augmentation experiments were conducted for both PMoC [31] and Triple-CFN [32]. These experiments involved expanding the data volume of the Bongard-Logo dataset using data augmentation techniques, specifically including rotation and flipping. It should be emphasized that these augmentation techniques do not explicitly alter the diversity of concepts or the proportion between concepts inherent in the Bongard-Logo dataset [31]. Rather, the aim of these experiments was to investigate the potential of the models with a larger dataset. Since PMoC posits that Transformer-based visual processing models require substantial data as a foundation, a point that is also echoed by ViT [17] and RS-Tran [28]. Furthermore, both PMoC and Triple-CFN agree that the utilization of these data augmentation methods serves to enrich the expressions of concepts at the pixel level, thereby challenging the network’s abilities to summarize, derive, accommodate, and invert concepts [31, 32]. To verify the effectiveness of D4C on the augmented dataset, we conducted experiments, and the results are presented in Table IV.
TABLE IV: Reasoning Accuracies of D4C on Augmented Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
PMoC 94.5 92.6 98.0 78.3 76.5
Triple-CFN 94.9 93.0 99.2 80.8 79.1
PMoC+D4C 96.0 95.2 99.2 84.8 82.8
Triple-CFN+D4C 97.7 94.8 99.6 82.8 83.1
The results presented in Table IV demonstrate that the integration of D4C into PMoC, which is under a probabilistic labeling scheme, has led to a significant enhancement in performance. The increase in data volume has enabled D4C to further narrow the gap between PMoC and Triple-CFN.
VIII-C Experiment on PGM
In this paper, we conducted experiments related to Lico-Net on PGM. And the results are recorded in Tabel V.
TABLE V: Reasoning Accuracies of models on PGM.
Model Test Accuracy(%)
SAVIR-T [24] 91.2
SCL [23, 24] 88.9
MRNet [21] 94.5
RS-CNN[28] 82.8
RS-TRAN[28] 97.5
Triple-CFN[32] 97.8
Triple-CFN+Re-space layer[32] 98.2
Triple-CFN+Re-space layer+D2C 98.3(98.25)
Lico-Net 98.5(98.52)
Lico-Net+D2C 98.5(98.53)
D4C (with Lico-Net as the backbone) 98.7(98.70)
D4C glove-train(with Lico-Net as the backbone) 98.8(98.82)
RS-Tran+Tranclip[28] 99.0
Meta Triple-CFN[32] 98.4
Meta Triple-CFN+Re-space layer[32] 99.3
Lico-Net+D3C- 99.3(99.30)
Lico-Net+D3C 99.4(99.35)
Lico-Net+D3C-cos 99.4(99.38)
From the experimental results presented in Table V, several notable phenomena emerge. Triple-CFN[32] positions progressive patterns and minimal reasoning units at the same logical level, leading to a design flaw due to their mismatched logic. Despite this, Triple-CFN remains an impressive model. By leveraging the RPM problem’s accurate inductive bias in its design, Lico-Net has surpassed Triple-CFN in RPM challenges.
The combinations of Lico-Net+D2C and Triple-CFN+D2C showed minimal improvement, limited to reasoning accuracy percentile. This was expected, as RPM and Bongard-Logo differ. RPM cases share a concept space and attribute recognition, rendering D2C, designed for concept interference and erroneous coupling issues, ineffective for RPM. The combination of D4C and D4C glove-train notably enhanced Lico-Net’s performance. D4C, our core contribution, showed promising results on both RPM and Bongard-Logo problems, highlighting the significance of this work.
The combination of Lico-Net and D3C is noteworthy. To validate D3C method, we applied both D3C and D3C-cos to Lico-Net. Through experimentation, we found that D3C and D3C-cos brought improvements to Lico-Net. To demonstrate the effectiveness of D3C’s approach to describing abstract concepts through distributions, this paper conducted ablation experiments using “Lico-Net+D3C-”. Unlike “Lico-Net+D3C”, “Lico-Net+D3C-” encodes Meta data into vectors. In essence, the Lico-Net+D3C- can be interpreted as fixing the variance of the distributions established for all concepts in Lico-Net+D3C-cos to 0.
TABLE VI: Progressive Pattern Reasoning Accuracies and of models on PGM.
Accuracy(%)
Model shape line answer
RS-Tran+Tranclip[28] - - 99.0
Meta Triple-CFN[32] 99.5 99.9 98.4
Meta Triple-CFN+Re-space layer[32] 99.7 99.9 99.3
Lico-Net+D3C- 99.7 99.9 99.3
Lico-Net+D3C 99.7 99.9 99.4
Lico-Net+D3C-cos 99.7 99.9 99.4
Table VI records D3C’s reasoning accuracies for progressive patterns, emphasizing Lico-Net+D3C(-cos)’s interpretability. Due to resource constraints, this paper did not fully explore the application of Sinkhorn-based D3C to Lico-Net. We stopped training once results illustrated our point, given the significant computational demands even with four A100s. In terms of results, D3C-cos can serve as a stand of D3C. Despite impressive reasoning accuracy and interpretability with Lico-Net+D3C(-cos), the improvement over previous works using Meta data as auxiliary labels isn’t significant. This might be because the inclusion of Meta data lowers the RPM problem’s reasoning difficulty. More supervisory signals typically enhance the intelligent performance of the system.
Lico-Net+D3C still shows distinct advantages over other Meta data-based methods. Specifically, Tranclip[28] utilizes Meta data to pre-train a set of warm-start parameters for RS-Tran[28], achieving promising reasoning accuracy for RS-Tran. However, Tranclip fails to simultaneously endow RS-Tran with both the the interpretability of progressive patterns and high reasoning accuracy. Meta Triple-CFN[32] also relies on a specially designed Re-space Layer[32] to enhance both its interpretability and reasoning accuracy. Similarly, the Re-Space Layer requires a fully trained Meta Triple-CFN to provide warm-start parameters [32]. In contrast, Lico-Net+D3C operates effectively without these extras, showing a more direct, coherent and self-consistent logic. This paper argues that this is not an insignificant contribution, as warm-start behavior may introduce unexpected instabilities, which are not a concern for Lico-Net and D3C. For instance, Tranclip requires early stopping when using Meta data for training RS-Tran’s warm-start parameters[28], and neglecting this can hinder RS-Tran’s performance. Similarly, The Re-space Layer [32] relies on a well-trained Meta Triple-CFN to provide warm-start parameters [32], a process that demands considerable training time and epochs, with inadequate training posing risks. Both methods require manual control, whereas Lico-Net+D3C mitigates these issues.
IX Conclusion
D4C, building on D2C and D3C principles, explores the roles of positive and negative examples in abstract reasoning from a distribution perspective. Positive examples depict the mean for solution distribution, while negative ones limit variance. This seems to reflect a despairing fact: the upper limit of accuracy achieved by the discriminator may be constrained by the quality of positive and negative examples. Therefore, D4C emerged to address this phenomenon. D4C adopts an adversarial approach, where the generator creates challenging negative examples and the discriminator refines its skills through defense. This iterative process enhances both components, resulting in remarkable solutions for abstract reasoning tasks. This paper does not advocate for the interpretability of the samples generated by the D4C generator; they are solely used to improve the performance of the discriminator. For future, the D4C method presents a unique perspective and approach for abstract reasoning problems and even other discriminative tasks. Its exploration of the identities of positive and negative examples may provide insights for the development of other works.